A person in high heels has the person's entire weight momentarily placed on one heel. Calculate the pressure exerted on the floor by the heel if it has an area of 1.71 cm2 and the person
Answer #1

The pressure is P=mg/A=(50.9)(9.8)/1.71*10-4=2.86 MPa